OpenStudy (anonymous):

by expressing y-4x-x^2+8 in form y=a(x-p)^2+q, where p and q are constant, state a) the equation of the axis of symmetry b) the coordinates of the turning point on the curve

OpenStudy (anonymous):

can u express y=4x-x^2+8 in the form y=a(x-p)^2+q,

OpenStudy (anonymous):

@sauravshakya nope

OpenStudy (anonymous):

y=4x-x^2+8 y=-(x^2-4x)+8 y=-(x^2-2*x*2+2^2-2^2)+8 y=-(x-2)^2+8+8 y=-(x-2)^2+16

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

hint: for y=(x-h)^2+k Co-ordinates of turning point===> (h,k) ANd axis of symmetry is x=h